In 1.31 grams of a chloride of titanium , there is 0.528 grams of titanium. What is the emperical formula of this titanium chloride?
1
Expert's answer
2020-06-17T05:56:10-0400

Solution.

Titanium chloride formula: TiCln.

n(Ti) = 0,528/48 = 0,011 mol.

X 0,011

TiCln = Ti + nCl

1 1

n(TiCln) = 0,011 mol.

M(TiCln) = m/n = 1,31/0,011 = 119 g/mol.

M(nCl) = 119 - 48 = 71 g/mol.

M(Cl) = 35,5 g/mol.

n = 2.

Formula: TiCl2.

Answer: TiCl2